Grade 12th passElectrostatics

The effective capacitance of two capacitors in parallel combination is 25 times the effective capacitance when they are in series. The ratio of C2/C1 is
  1. 3/2
  2. 4/3
  3. 5/3
  4. 25/6

Yogiraj
10 Years agoGrade 12th pass

1 Answer

Vikas TU
10 Years ago
(C1 + C2)  =25(C1C2/(C1+C2))
(C1 + C2)^2 = 25C1C2
(C1 + C2 + 5root(C1C2)) (C1 +  C2 – 5root(C1C2)) = 0
(C1 + C2 - 5root(C1C2))  = 0
C1 + C2 = 5root(C1C2)
Divide the whole eqn. by C1 we get,
1 + C2/C1 = 5root(C2/C1)
let C2/C1 = x
after putting u will get a quadratic solve it and get vale value of x .i..e. C2/C1